A company has financed its total asset with 38% debt. What would the firms equity multiplier be?
Answer #1

TOTAL ASSETS = 100

DEBT = 38, SO EQUITY = TOTAL ASSETS - DEBT = 100 -38 = 62

EQUITY MULTIPLIER = TOTAL ASSETS/EQUITY = 100/62 = 1.6129 (rounded to 4 decimals]

one more way to calculate is :

EQUITY MULTIPLIER = 1+ (DEBT/EQUITY) = 1 + (38/62) = 1.6129 (rounded to 4 decimals) (Thumbs up please)